OpenStudy (anonymous):

True or false. cos(a/2)=sqrt(1-cos(a)/2) for all vaules of a.

OpenStudy (anonymous):

seems unlikely, since cosine can be negative

OpenStudy (anonymous):

actually now that i look, the whole thing is wrong \[\cos(\frac{\alpha}{2})=\pm\sqrt{\frac{1+\cos(\alpha)}{2}}\]

OpenStudy (anonymous):

SO THEN IT WOULD BE FALSE?

OpenStudy (anonymous):

definitely

OpenStudy (anonymous):

THANKS!